OpenStudy (anonymous):

Medal for helping!!! After how many months will Rita pay one hundred dollars in simple annual interest on a credit card that charges twelve percent on her five hundred dollars balance? A. 2 months B. 4 months C. fourteen months D. twenty months

OpenStudy (mathmale):

Let's concentrate on HOW we calculate the length of time involved. The formula for simple interest is i = p*r*t, where p is the principal, r the interest rate (as a decimal fraction), and t is the time (in years). We want t, so why not solve i = p*r*t for t? Then substitute p=$500, i=$100, r=.12 and calculate t. That will be n years. Convert that to months.

OpenStudy (anonymous):

Genius!

OpenStudy (anonymous):

It gives me 6000

OpenStudy (mathmale):

Dear CashBoy, please try calculating t again: t = i/(pr). 6000 months would be 500 years, and you know that's not the correct answer.

OpenStudy (anonymous):

True

OpenStudy (mathmale):

So, CashBoy, t = i/(pr) becomes t = ($100)/($500*0.12) = ? (in years). Convert this reslt to the number of months.